Oscar, Inc. issued a $500,000, 5.5%, 10-year bond on January 1, 2016. The market rate of interest at issuance was 6%. The bonds make semi-annual interest payments on July 1st and January 1st. The corporation’s fiscal year ends on December 31. The corporation accounts for the bonds using the effective-interest method

a) Calculate the price of the bond at issuance on January 1, 2016

b) Prepare any journal entry the company records on December 31, 2016
Answer #1

a)Semiannual interest =Par value *coupon rate *n/12

                        = 500000 * .055 *6/12                [2 semiannual period in a year comprising of 6 months each]

                         = 13750

semiannual period = 10*2 =20

semiannual yield (market rate) = 6*6/12 = 3%

Price of bond =[PVA3%,20*semiannual interest ] +[PVF3%,20*Par value]

                =[14.87747 * 13750 ] +[.55368 * 500000]

                = 204565.21+ 276840

               = 481405 rounded

#Find present value factor from present value table at 3% for 20 periods or using the formula 1/(1+i)^n

#Find present value annuity factor from present value annuity table at 3% for 20 periods

b)

Date Interest expense [A] Interest paid [B] Discount amortization [A-B] Carrying value at end
1 Jan 481405
1July 481405*3%= 14442 13750 692 481405+692= 482097
31 Dec 482097*3%=14463 13750 713 482097+713= 482810
Date Account title Debit credit
December 31, 2016 Interest expense 14463
Discount on bond payable 713
Interest payable 13750
[being interest accrued for semiannual period ending December 31 to be paid on 1Jan 2017